# On Arratia’s coupling and the Dirichlet law for the factors of a random integer

### v1

@@ 9. A coupling for $\mathbf{D}_{x}$
Fix $k\geqslant 2$, $\bm{\alpha}\in\Delta^{k-1}$ with $\alpha_{i}>0$ for all $i$, and $f\in\mathcal{F}_{k}(\bm{\alpha})$. In addition, let $\mathbf{D}_{x}$ be a random variable satisfying (1.5), and let $\mathbf{Z}$ be a $\Delta^{k-1}$-valued random variable distributed according to $\text{Dir}(\bm{\alpha})$. We will construct an appropriate coupling between $\mathbf{D}_{x}$ and $\mathbf{Z}$. This will build on the coupling given in Section 2. To state our result, we also introduce the $\Delta^{k-1}$-valued random variable
@@
$R_{\mathrm{NT}}\coloneqq 3\cdot\frac{\log(x/N_{x})+\log s(N_{x})}{\log x}\quad\text{and}\quad R_{\mathrm{PD}}\coloneqq 2\cdot\frac{\Theta_{x}}{\log x}.$
@@
Indeed, assume that $\mathcal{E}\cap\mathcal{E}_{\mathrm{NT}}(A)\cap\mathcal{E}_{\mathrm{PD}}(A)$ occurs, and we also have $\bm{\delta}_{x}\in A$. We must show that $\mathbf{Z}\in A$. Assume for contradiction that $\mathbf{Z}\notin A$.
@@
On the other hand, since $\mathbf{Z}\in\Delta^{k-1}\setminus A$ and $B(\mathbf{Z},R_{\mathrm{PD}})\cap\partial A=\emptyset$, we must have
@@
$\mathcal{E}_{\mathrm{NT}}(\Delta^{k-1}_{\mathbf{u}})^{c}\subseteq\bigcup_{\begin{subarray}{c}i<k\\ u_{i}\neq 1\end{subarray}}\mathcal{B}_{\mathrm{NT}}^{(i)}(u_{i}),\quad\text{where}\quad\mathcal{B}_{\mathrm{NT}}^{(i)}(u)\coloneqq\Big{\{}\big{|}\log D_{x,i}-u\log x\big{|}\leqslant 3\log(x\cdot s(N_{x})/N_{x})\Big{\}},$
@@
Let $n^{\flat}=dm$ and $s(n)=b$. If $n^{\flat}\in[\frac{x}{z},\frac{2x}{z}]$ and $d\in[x^{u}z^{-4},x^{u}z^{3}]$, then $m\in[x^{1-u}z^{-2},2x^{1-u}z^{3}]$. Hence, we find that
$S(z)$ $\leqslant\mathop{\sum\sum}_{\begin{subarray}{c}dm\in[\frac{x}{z},\frac{2x}{z}]\\ d\in[x^{u}z^{-4},x^{u}z^{3}]\\ m\in[x^{1-u}z^{-2},2x^{1-u}z^{3}]\end{subarray}}\alpha^{\omega(d)}(1-\alpha)^{\omega(m)}\sum_{\begin{subarray}{c}b\leqslant z\\ b\ \text{square-full}\end{subarray}}1$
(10.3) $\ll\sqrt{z}\mathop{\sum\sum}_{\begin{subarray}{c}dm\in[\frac{x}{z},\frac{2x}{z}]\\ d\in[x^{u}z^{-4},x^{u}z^{3}]\\ m\in[x^{1-u}z^{-2},2x^{1-u}z^{3}]\end{subarray}}\alpha^{\omega(d)}(1-\alpha)^{\omega(m)}=\sqrt{z}\cdot(S_{1}+S_{2}),$
@@
$S_{2}$ $\leqslant\sum_{\begin{subarray}{c}m\leqslant\sqrt{2x/z}\\ m\in[x^{1-u}z^{-3},2x^{1-u}z^{2}]\end{subarray}}(1-\alpha)^{\omega(m)}\sum_{d\leqslant\frac{2x}{mz}}\alpha^{\omega(d)}$
$\ll\sum_{\begin{subarray}{c}m\leqslant\sqrt{2x/z}\\ m\in[x^{1-u}z^{-3},2x^{1-u}z^{2}]\end{subarray}}(1-\alpha)^{\omega(m)}\cdot\frac{x(\log x)^{\alpha-1}}{mz}$
@@
Uniformly for $u,\delta,\alpha\in[0,1]$, we have
@@
The lemma holds trivially when $\alpha\in\{0,1\}$, so let us assume that $\alpha\in(0,1)$. Note that both sides of the claimed inequality are invariant under the change of variables $(\alpha,u)\to(1-\alpha,1-u)$. Hence, we may assume without loss of generality that $u\in[0,1/2]$. In addition, observe that
@@
Hence the lemma is trivially true if $\delta\in[1/4,1]$.
@@
Let us now show Lemma 9.4. For the simplicity of notation, let us fix $i\in\{1,2,\dots,k\}$ and $u\in(0,1)$, and let us set $\alpha=\alpha_{i}$ and
$\Delta=\frac{1}{(1+u\log x)^{1-\alpha}(1+(1-u)\log x)^{\alpha}}.$
@@
On the other hand, if $\frac{200c}{\log x}<|Z_{i}-u|\leqslant\frac{2\Theta_{x}}{\log x}\leqslant\frac{20c+2\Theta_{x}^{\prime}}{\log x}$, then we must have $\Theta_{x}^{\prime}>90c$. In particular, there exists $m\in\mathbb{Z}$ such that $2^{m}>40c$ and $\Theta_{x}^{\prime}\in(2^{m},2^{m+1}]$, whence $|Z_{i}-u|\leqslant 2^{m+2}/\log x$. We thus conclude that
@@
$\begin{split}\mathbb{P}\bigg{[}\big{|}Z_{i}-u\big{|}\leqslant\frac{20c+4\kappa}{\log x},\ \Theta_{x}^{\prime}>\kappa\bigg{]}&\leqslant\sum_{\begin{subarray}{c}m\geqslant 0\\ 2^{m}\leqslant 0.1\log x\end{subarray}}\frac{250c^{2}}{\kappa^{2}2^{3m}}\mathbb{E}\Big{[}\mathbbm{1}_{|Z_{i}-u|\leqslant\frac{20c+4\kappa}{\log x}}\cdot\mathbbm{1}_{G(2^{m})\geqslant 4}\cdot G(2^{m})^{2}\Big{]}\\ &\ +\sum_{m<0}\frac{250c^{2}2^{m}}{\kappa^{2}}\mathbb{E}\Big{[}\mathbbm{1}_{|Z_{i}-u|\leqslant\frac{20c+4\kappa}{\log x}}\cdot\mathbbm{1}_{G(2^{m})\geqslant 4}\cdot G(2^{m})^{2}\Big{]}.\end{split}$
@@
where $G_{A}(\lambda)=\sum_{j\geqslant 1}\mathbbm{1}_{A_{j}\log x\in(\lambda S,2\lambda S]}$ and $G_{B}(\lambda)=\sum_{j\geqslant 1}\mathbbm{1}_{B_{j}\log x\in(\lambda S,2\lambda S]}$. We know $G_{A}(\lambda)+G_{B}(\lambda)\geqslant 4$. Hence, if $G_{A}(\lambda)\geqslant G_{B}(\lambda)$, then we must have $G_{A}(\lambda)\geqslant 2$, whence $G_{A}(\lambda)\leqslant G_{A}(\lambda)^{2}/2$. So, we find that
@@
$\begin{split}E_{A}&=\iint\limits_{2a_{1}>a_{2}>a_{1}>0}\mathbb{E}\bigg{[}\mathbbm{1}_{|\frac{a_{1}+a_{2}+S_{A}}{a_{1}+a_{2}+S}-u|\leqslant\frac{\mu}{\log x}}\prod_{j=1}^{2}\mathbbm{1}_{a_{j}\log x\in(\lambda(a_{1}+a_{2}+S),2\lambda(a_{1}+a_{2}+S)]}\bigg{]}\frac{e^{-a_{1}-a_{2}}}{a_{1}a_{2}}\,\mathrm{d}a_{1}\,\mathrm{d}a_{2}\\ &=\frac{1}{\Gamma(\alpha)\Gamma(1-\alpha)}\iiiint\limits_{\begin{subarray}{c}2a_{1}>a_{2}>a_{1}>0,\ s_{1},s_{2}>0\\ |\frac{a_{1}+a_{2}+s_{1}}{a_{1}+a_{2}+s_{1}+s_{2}}-u|\leqslant\frac{\mu}{\log x}\\ a_{j}\log x\in(\lambda(a_{1}+a_{2}+s_{1}+s_{2}),2\lambda(a_{1}+a_{2}+s_{1}+s_{2})]\ (j=1,2)\end{subarray}}\frac{e^{-a_{1}-a_{2}-s_{1}-s_{2}}}{a_{1}a_{2}s_{1}^{1-\alpha}s_{2}^{\alpha}}\,\mathrm{d}a_{1}\,\mathrm{d}a_{2}\,\mathrm{d}s_{1}\,\mathrm{d}s_{2},\end{split}$
where we used that the $a_{j}$’s lie in the same dyadic interval to deduce that $a_{2}<2a_{1}$. We make the change of variables $t=s_{1}/(s_{1}+s_{2})$ and $s=s_{1}+s_{2}$. Since $\lambda/\log x\leqslant 0.1$ and $a_{1}<a_{2}<2a_{1}$, the conditions $a_{j}\log x\in\lambda(a_{1}+a_{2}+s),2\lambda(a_{1}+a_{2}+s)]$ for $j=1,2$ imply that $a_{j}\log x\in(\lambda s,5\lambda s]$. Knowing also that $|\frac{a_{1}+a_{2}+s_{1}}{a_{1}+a_{2}+s}-u|\leqslant\frac{\mu}{\log x}$, we find $|t-u|=|\frac{s_{1}}{s_{1}+s_{2}}-u|\leqslant\frac{10\lambda+\mu}{\log x}$. Finally, we use Euler’s reflection formula to write $\Gamma(\alpha)\Gamma(1-\alpha)=\pi/\sin(\pi\alpha)$. We conclude that
@@
For every fixed value of $s$, the integral over $a_{1}$ and $a_{2}$ is $\leqslant(\log 5)^{2}$, since $\int_{w}^{5w}\frac{e^{-a}}{a}\,\mathrm{d}\leqslant\log 5$ for any $w>0$. We also have $\int_{0}^{\infty}e^{-s}\,\mathrm{d}s=1$. We thus conclude that
@@ Appendix A Tools from Number Theory
$\sum_{p^{k}\leqslant x}\frac{1}{kp^{k}}=\int_{2^{-}}^{x}\frac{\mathrm{d}\psi(t)}{t\log t}=\log\log x+c+O\!\left(\frac{1}{(\log x)^{3}}\right)$
@@
The first bound follows readily by Theorem 14.2 in [12]. Let us now prove the second one.
@@ B.2. Poisson Processes
is absolutely convergent almost surely if and only if
@@ Appendix C Tools from Topology
Let $(X,\mathcal{O})$ be a topological space, where $\mathcal{O}$ is the set of open sets of $X$, and let $A\subseteq X$.
@@
    We define the boundary of $A$ to equal $\partial A\coloneqq\overline{A}\setminus\text{int}(X\setminus A)=\overline{A}\cap\overline{X\setminus A}$.
@@
    We say that $A$ is disconnected if there exist two open sets $O_{1},O_{2}$ such that $A=(A\cap O_{1})\sqcup(A\cap O_{2})$. We say that $A$ is connected if it is not disconnected.
@@
Let $X$ be a topological space, let $A\subseteq X$, and let $B$ be a connected subset of $\mathbb{R}^{k}$ such that $B\cap\partial A=\emptyset$. Then either $B\subseteq A$ or $B\subseteq X\setminus A$.
@@
Since $\partial A=\overline{A}\cap\overline{A^{c}}$, we have that $\partial A,\text{int}(A)$ and $\text{int}(X\setminus A)$ are disjoint sets partitioning $X$. By our assumption that $B\cap\partial A=\emptyset$, we must thus have $B\subseteq\text{int}(A)\sqcup\text{int}(X\setminus A)$. By our assumption that $B$ is connected, we must then have that either $B\subseteq\text{int}(A)\subseteq A$ or that $B\subseteq\text{int}(X\setminus A)\subseteq X\setminus A$. This completes the proof. ∎

### v2

@@ 9. A coupling for $\mathbf{D}_{x}$
Fix $k\geqslant 2$, $\bm{\alpha}\in\Delta^{k-1}$ with $\alpha_{i}>0$ for all $i$, and $f\in\mathcal{F}_{k}(\bm{\alpha})$. In addition, let $\mathbf{D}_{x}$ be a random variable satisfying (1.5), and let $\mathbf{Z}$ be a $\Delta^{k-1}$-valued random variable distributed according to $\text{Dir}(\bm{\alpha})$. We will construct an appropriate coupling between $\mathbf{D}_{x}$ and $\mathbf{Z}$. This will build on the coupling given in Section 2. To state our result, we also introduce the random vectors
@@
$R_{\mathrm{NT}}\coloneqq 3\cdot\frac{\log(x/N_{x}^{\flat})}{\log x}\quad\text{and}\quad R_{\mathrm{PD}}\coloneqq 2\cdot\frac{\Theta_{x}}{\log x}.$
@@
Indeed, assume that $\mathcal{E}\cap\mathcal{E}_{\mathrm{NT}}(A)\cap\mathcal{E}_{\mathrm{PD}}(A)$ occurs, and we also have $\bm{\delta}_{x}\in A$. Aiming for a contradiction, we also assume that $\mathbf{Z}\notin A$.
@@
And by assumption of $\mathcal{E}_{\mathrm{PD}}(A)$, we have $\mathbf{Z}\in\Delta^{k-1}\setminus A$ and $B(\mathbf{Z},R_{\mathrm{PD}})\cap\partial A=\emptyset$, which implies
@@
$\mathcal{E}_{\mathrm{NT}}(\Delta^{k-1}_{\mathbf{u}})^{c}\subseteq\bigcup_{\begin{subarray}{c}i<k\\ u_{i}\neq 1\end{subarray}}\mathcal{B}_{\mathrm{NT}}^{(i)}(u_{i}),\quad\text{where}\quad\mathcal{B}_{\mathrm{NT}}^{(i)}(u)\coloneqq\Big{\{}\big{|}\log D_{x,i}-u\log x\big{|}\leqslant 3\log(x/N_{x}^{\flat})\Big{\}},$
@@
Let $n^{\flat}=dm$ and $s(n)=b$. If $n^{\flat}\in[\frac{x}{z},\frac{2x}{z}]$ and $d\in[x^{u}z^{-4},x^{u}z^{3}]$, then $m\in[x^{1-u}z^{-4},2x^{1-u}z^{3}]$. Hence, we find that
$S(z)$ $\leqslant\mathop{\sum\sum}_{\begin{subarray}{c}dm\in[\frac{x}{z},\frac{2x}{z}]\\ d\in[x^{u}z^{-4},x^{u}z^{3}]\\ m\in[x^{1-u}z^{-4},2x^{1-u}z^{3}]\end{subarray}}\alpha^{\omega(d)}(1-\alpha)^{\omega(m)}\sum_{\begin{subarray}{c}b\leqslant z\\ b\ \text{square-full}\end{subarray}}1$
(10.3) $\ll\sqrt{z}\mathop{\sum\sum}_{\begin{subarray}{c}dm\in[\frac{x}{z},\frac{2x}{z}]\\ d\in[x^{u}z^{-4},x^{u}z^{3}]\\ m\in[x^{1-u}z^{-4},2x^{1-u}z^{3}]\end{subarray}}\alpha^{\omega(d)}(1-\alpha)^{\omega(m)}=\sqrt{z}\cdot(S_{1}+S_{2}),$
@@
$S_{2}$ $\leqslant\sum_{\begin{subarray}{c}m\leqslant\sqrt{2x/z}\\ m\in[x^{1-u}z^{-4},2x^{1-u}z^{3}]\end{subarray}}(1-\alpha)^{\omega(m)}\sum_{d\leqslant\frac{2x}{mz}}\alpha^{\omega(d)}$
$\ll\sum_{\begin{subarray}{c}m\leqslant\sqrt{2x/z}\\ m\in[x^{1-u}z^{-4},2x^{1-u}z^{3}]\end{subarray}}(1-\alpha)^{\omega(m)}\cdot\frac{x(\log x)^{\alpha-1}}{mz}$
@@
Uniformly for $u\in[0,1]$, $\alpha\in(0,1)$ and $\delta>0$, we have
@@
Note that both sides of the claimed inequality are invariant under the change of variables $(\alpha,u)\mapsto(1-\alpha,1-u)$. Hence, we may assume without loss of generality that $u\in[0,1/2]$. In addition, observe that
@@
Hence the lemma is trivially true if $\delta>1/4$.
@@
Let us now show Lemma 9.4. For simplicity of notation, let us fix $i\in\{1,2,\dots,k\}$ and $u\in(0,1)$, and let us set $\alpha\coloneqq\alpha_{i}$ and
$\Delta\coloneqq\frac{1}{(1+u\log x)^{1-\alpha}(1+(1-u)\log x)^{\alpha}}.$
@@
On the other hand, if $\frac{200c}{\log x}<|Z_{i}-u|\leqslant\frac{2\Theta_{x}}{\log x}\leqslant\frac{20c+2\Theta_{x}^{\prime}}{\log x}$, then we must have $\Theta_{x}^{\prime}>90c$. In particular, there exists $m\in\mathbb{Z}$ such that $2^{m}>40c$ and $\Theta_{x}^{\prime}\in(2^{m},2^{m+1}]$, whence $|Z_{i}-u|\leqslant\frac{20c+2^{m+2}}{\log x}$. We thus conclude that
@@
$\begin{split}\mathbb{P}\bigg{[}\big{|}Z_{i}-u\big{|}\leqslant\frac{20c+4\kappa}{\log x},\ \Theta_{x}^{\prime}>\kappa\bigg{]}&\leqslant\sum_{\begin{subarray}{c}m\geqslant 0\\ 2^{m}\leqslant 0.1\log x\end{subarray}}\frac{100c^{2}}{\kappa^{2}2^{3m}}\mathbb{E}\Big{[}\mathbbm{1}_{|Z_{i}-u|\leqslant\frac{20c+4\kappa}{\log x}}\cdot\mathbbm{1}_{G(2^{m})\geqslant 4}\cdot G(2^{m})^{2}\Big{]}\\ &\ +\sum_{m<0}\frac{100c^{2}2^{m}}{\kappa^{2}}\mathbb{E}\Big{[}\mathbbm{1}_{|Z_{i}-u|\leqslant\frac{20c+4\kappa}{\log x}}\cdot\mathbbm{1}_{G(2^{m})\geqslant 4}\cdot G(2^{m})^{2}\Big{]}.\end{split}$
@@
where $G_{A}(\lambda)=\sum_{j\geqslant 1}\mathbbm{1}_{A_{j}\log x\in(\lambda S,2\lambda S]}$ and $G_{B}(\lambda)=\sum_{j\geqslant 1}\mathbbm{1}_{B_{j}\log x\in(\lambda S,2\lambda S]}$. We know $G_{A}(\lambda)+G_{B}(\lambda)\geqslant 4$. Hence, if $G_{A}(\lambda)\geqslant G_{B}(\lambda)$, then we must have $G_{A}(\lambda)\geqslant 2$ and $G_{A}(\lambda)\leqslant G_{A}(\lambda)^{2}/2$, whence $G_{A}(\lambda)\leqslant 2\mathop{\sum\sum}_{k>j\geqslant 1}\mathbbm{1}_{A_{k}\log x,\,A_{j}\log x\in(\lambda S,2\lambda S]}$. So, we find that
@@
$\begin{split}E_{A}&=\alpha^{2}\iint\limits_{2a_{1}>a_{2}>a_{1}>0}\mathbb{E}\bigg{[}\mathbbm{1}_{|\frac{a_{1}+a_{2}+S_{A}}{a_{1}+a_{2}+S}-u|\leqslant\frac{\mu}{\log x}}\prod_{j=1}^{2}\mathbbm{1}_{a_{j}\log x\in(\lambda(a_{1}+a_{2}+S),2\lambda(a_{1}+a_{2}+S)]}\bigg{]}\frac{e^{-a_{1}-a_{2}}}{a_{1}a_{2}}\,\mathrm{d}a_{1}\,\mathrm{d}a_{2}\\ &=\frac{\alpha^{2}}{\Gamma(\alpha)\Gamma(1-\alpha)}\iiiint\limits_{\begin{subarray}{c}2a_{1}>a_{2}>a_{1}>0,\ s_{1},s_{2}>0\\ |\frac{a_{1}+a_{2}+s_{1}}{a_{1}+a_{2}+s_{1}+s_{2}}-u|\leqslant\frac{\mu}{\log x}\\ a_{j}\log x\in(\lambda(a_{1}+a_{2}+s_{1}+s_{2}),2\lambda(a_{1}+a_{2}+s_{1}+s_{2})]\ (j=1,2)\end{subarray}}\frac{e^{-a_{1}-a_{2}-s_{1}-s_{2}}}{a_{1}a_{2}s_{1}^{1-\alpha}s_{2}^{\alpha}}\,\mathrm{d}a_{1}\,\mathrm{d}a_{2}\,\mathrm{d}s_{1}\,\mathrm{d}s_{2},\end{split}$
where we used that the $a_{j}$’s lie in the same dyadic interval to deduce that $a_{2}<2a_{1}$. We make the change of variables $t=s_{1}/(s_{1}+s_{2})$ and $s=s_{1}+s_{2}$. Since $\lambda/\log x\leqslant 0.1$ and $a_{1}<a_{2}<2a_{1}$, the conditions $a_{j}\log x\in(\lambda(a_{1}+a_{2}+s),2\lambda(a_{1}+a_{2}+s)]$ for $j=1,2$ imply that $a_{j}\log x\in(\lambda s,5\lambda s]$. Knowing also that $|\frac{a_{1}+a_{2}+s_{1}}{a_{1}+a_{2}+s}-u|\leqslant\frac{\mu}{\log x}$, we find $|t-u|=|\frac{s_{1}}{s_{1}+s_{2}}-u|\leqslant\frac{10\lambda+\mu}{\log x}$. Finally, we use Euler’s reflection formula to write $\Gamma(\alpha)\Gamma(1-\alpha)=\pi/\sin(\pi\alpha)$. We conclude that
@@
For every fixed value of $s$, the integral over $a_{1}$ and $a_{2}$ is $\leqslant(\log 5)^{2}$, since $\int_{w}^{5w}\frac{e^{-a}}{a}\,\mathrm{d}a\leqslant\log 5$ for any $w>0$. We also have $\int_{0}^{\infty}e^{-s}\,\mathrm{d}s=1$. We thus conclude that
@@ Appendix A Tools from Number Theory
$\sum_{p^{k}\leqslant x}\frac{1}{kp^{k}}=\int_{2^{-}}^{x}\frac{\mathrm{d}\psi(t)}{t\log t}=\log_{2}x+c+O\!\left(\frac{1}{(\log x)^{3}}\right)$
@@
The first bound follows readily with Theorem 14.2 in [12]. Let us now prove the second one.
@@ B.2. Poisson Processes
is absolutely convergent almost surely if, and only if,
@@ Appendix C Tools from Topology
Let $(X,\mathcal{O})$ be a topological space, and let $A\subseteq X$.
@@
    We define the boundary of $A$ to equal $\partial A\coloneqq\overline{A}\setminus\text{int}(A)$.
@@
    We say that $A$ is disconnected if there exist two disjoint open sets $O_{1}$ and $O_{2}$ such that $A\subseteq O_{1}\sqcup O_{2}$ with $A\cap O_{1}$ and $A\cap O_{2}$ being both non-empty. We say that $A$ is connected if it is not disconnected.
@@
Let $X$ be a topological space and let $A,B\subseteq X$ such that $B\cap\partial A=\emptyset$ and $B$ is connected. Then either $B\subseteq A$ or $B\subseteq X\setminus A$.
@@
From the definition of closure and boundary above, the sets $\partial A,\text{int}(A)$ and $\text{int}(X\setminus A)$ form a partition of $X$. By our assumption that $B\cap\partial A=\emptyset$, we must thus have $B\subseteq\text{int}(A)\sqcup\text{int}(X\setminus A)$. By our assumption that $B$ is connected, we must then have that either $B\subseteq\text{int}(A)\subseteq A$ or that $B\subseteq\text{int}(X\setminus A)\subseteq X\setminus A$. This completes the proof. ∎